# Iterating $\mathsf{PP}$-packages without Choice: A Cohen symmetric seed and a localization framework

### v1

@@ 2 Preliminaries and black-box iteration technology
    Set-length and class-length versions of symmetric iterations (cf. [3, §2.1]).
@@
    Definitions and basic properties of countable-support symmetric iterations (cf. [3, Def. 3.1]).
@@
    A $\mathsf{DC}$-preservation theorem for the relevant support (cf. [3, Thm. 3.20]).
@@
    Generation and completeness/normality of limit filters in the relevant cofinality cases (cf. [3, Thm. 5.5]).
@@
    A characterization of hereditarily symmetric names by “small support” (cf. [3, Lem. 3.14]).
@@
The following facts from [13] drive our localization strategy:
Fact 2.3.
For any set $S$, $\mathsf{SVC}(S)$ implies $\mathsf{SVC}^{+}(\mathcal{P}(S))$.
Theorem 2.4 (Ryan–Smith Localization, [13]).
Assume $\mathsf{SVC}^{+}(T)$. Then $\mathsf{PP}$ is equivalent to the conjunction of $\mathsf{PP}\!\restriction\!T$ and $\mathsf{AC}_{\mathsf{WO}}$.
Our strategy is to construct a model satisfying $\mathsf{SVC}(S)$ (via a Cohen seed) and then force $\mathsf{PP}\!\restriction\!\mathcal{P}(S)$ and $\mathsf{AC}_{\mathsf{WO}}$ while preserving $\mathsf{SVC}(S)$.
@@ 3.1 The forcing, group action, and the countable-support filter
But $\pi_{i}\operatorname{Fix}(E_{i})\pi_{i}^{-1}=\operatorname{Fix}(\pi_{i}[E_{i}])$, and finite intersections of $\operatorname{Fix}(\cdot)$’s are $\operatorname{Fix}(\bigcup\cdot)$. Let $E:=\bigcup_{i<k}\pi_{i}[E_{i}]$ (countable). Then $\operatorname{Fix}(E)\leq H$. ∎
@@
Let $\langle H_{n}:n<\omega\rangle\subseteq\mathscr{F}$. By Lemma 3.4 pick countable $E_{n}$ with $\operatorname{Fix}(E_{n})\leq H_{n}$. Put $E=\bigcup_{n<\omega}E_{n}$ (countable), so $\operatorname{Fix}(E)=\bigcap_{n<\omega}\operatorname{Fix}(E_{n})\leq\bigcap_{n<\omega}H_{n}$, hence $\bigcap_{n<\omega}H_{n}\in\mathscr{F}$. ∎
@@ 3.2 CCC and dependent choice
Let $D_{n+1}:=\bigcup_{p\in D_{n}}D_{n+1,p}$. Then $D_{n+1}$ is countable (a countable union of countable sets) and is a maximal antichain below $p_{\ast}$: given any $r\leq p_{\ast}$, pick $p\in D_{n}$ compatible with $r$ (maximality of $D_{n}$), let $r^{\prime}\leq r,p$, and then pick $q\in D_{n+1,p}$ compatible with $r^{\prime}$ (maximality of $D_{n+1,p}$).
@@
a countable union of countable sets of ordinals, hence countable. Then every $\pi\in\operatorname{Fix}(E_{n})$ fixes each $p\in D_{n}$, fixes $\dot{A}$ and $\dot{R}$, and fixes each $\dot{a}_{n,p}$. Therefore $\pi\dot{a}_{n}=\dot{a}_{n}$, i.e. $\operatorname{Fix}(E_{n})\leq\operatorname{Stab}(\dot{a}_{n})$, so $\dot{a}_{n}$ is symmetric; hereditariness follows since the names appearing in $\dot{a}_{n}$ are among the $\dot{a}_{n,p}\in\operatorname{HS}$. Thus $\dot{a}_{n}\in\operatorname{HS}$ for all $n$.
Let $E:=\bigcup_{n<\omega}E_{n}$, again countable, and define the sequence name
@@ 3.3 Countable supports for $\operatorname{HS}$-names
Corollary 3.16 (Supported names live in the corresponding subextension).
Let $E\subseteq\omega_{1}$ and $\tau$ a $\mathbb{P}$-name with $\operatorname{Fix}(E)\leq\operatorname{Stab}(\tau)$. Then $\mathbbm{1}_{\mathbb{P}}\Vdash\tau=\tau\!\restriction E$. In particular, for any $V$-generic $G\subseteq\mathbb{P}$ with $G_{E}:=G\cap\mathbb{P}_{E}$,
$\tau^{G}=(\tau\!\restriction E)^{G_{E}}\in V[G_{E}].$
@@
We prove by induction on rank that $\mathbbm{1}_{\mathbb{P}}\Vdash\tau=\tau\!\restriction E$. Fix a formula of the form $\sigma\in\tau$” or $\sigma=\tau$” and use Lemma 3.15 to show that whenever $p$ decides such a statement about $\tau$, its restriction $p\!\restriction E$ decides it the same way. This yields that membership in $\tau^{G}$ is determined by $G_{E}$ and the recursive definition of $\tau\!\restriction E$, and the induction closes. ∎
@@ 3.4 The canonical Cohen set $A$ and failure of choice
Unlike the finite-support generalized Cohen model (which produces a Dedekind-finite $A$; cf. [11, §2.4]), the present countable-support filter allows many countable subsets of $A$: for each countable $E\subseteq\omega_{1}$ in $V$, the name $\dot{A}_{E}:=\{(\dot{c}_{\alpha},\mathbf{1}):\alpha\in E\}$ is hereditarily symmetric, so $A_{E}:=\{c_{\alpha}:\alpha\in E\}\in\mathcal{N}$.
@@ 3.5 A safe $\mathsf{SVC}$ seed in $\mathcal{N}$
Fix a nonempty $X\in\mathcal{N}$ and choose $\dot{X}\in\operatorname{HS}$ with $X=\dot{X}^{G}$.
By Lemma 3.13, fix a countable support $E\subseteq\omega_{1}$ for $\dot{X}$, i.e. $\operatorname{Fix}(E)\leq\operatorname{Stab}(\dot{X})$. Let $\dot{X}\!\restriction E$ be the restricted $\mathbb{P}_{E}$-name from Notation above. By Corollary 3.16,
$X=\dot{X}^{G}=(\dot{X}\!\restriction E)^{G_{E}}\in V[G_{E}],$
where $G_{E}:=G\cap\mathbb{P}_{E}$.
Since $V[G_{E}]\models\mathsf{ZFC}$, there is an ordinal $\eta$ and a surjection $h:\eta\twoheadrightarrow X$ in $V[G_{E}]$. Let $\dot{h}$ be a $\mathbb{P}_{E}$-name for $h$. Every permutation $\pi\in\operatorname{Sym}(\omega_{1})$ which fixes $E$ pointwise acts trivially on $\mathbb{P}_{E}$ and hence fixes $\dot{h}$, so $\dot{h}\in\operatorname{HS}$ (indeed, it has support $E$). Therefore $h=\dot{h}^{G}\in\mathcal{N}$.
Finally, since $S=A^{\omega}$ is nonempty in $\mathcal{N}$, the map $f:S\times\eta\to X$ defined by $f(s,\xi)=h(\xi)$ is a surjection in $\mathcal{N}$. This proves $\mathcal{N}\models\mathsf{SVC}(S)$. ∎
@@ 4 Reduction blueprint: global $\mathsf{PP}$ from localized $\mathsf{PP}$
For every set $S$, $\mathsf{SVC}(S)$ implies $\mathsf{SVC}^{+}(\mathcal{P}(S))$. [13, §2.1, Fact]
@@
[13, Proposition 3.17] (stated there with parameter “SS”; we apply it with $S:=T$)
@@
This is witnessed by Cohen’s first model; see [13, Proposition 3.22]. ∎
@@
This is witnessed by the Feferman-style model $N_{\aleph_{1}}$ of Truss; see [13, Proposition 3.23]. ∎
@@
The preceding propositions are included only to justify (conceptually) why the localization theorem is stated with both components present: the “choice” side $\mathsf{AC}_{\mathsf{WO}}$ and the “localized comparability” side $\mathsf{PP}\!\restriction T$. We will force both conjuncts $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ in Theorem 4.5. In Ryan–Smith’s analysis the corresponding reduction is sharp: in general one cannot drop the displayed conjuncts (nor substantially weaken the hypotheses) without losing the equivalence; see [13, Props. 3.22 and 3.23].

### v3

@@ 1 Introduction
Remark 1.1 (The axiom $\mathsf{AC}_{\mathsf{WO}}$).
Throughout this paper, $\mathsf{AC}_{\mathsf{WO}}$ denotes the axiom commonly called choice for well–orderable sets, in the sense of Howard–Rubin [7, Form 40]. That is:
Every well–orderable set of nonempty sets admits a choice function.
Equivalently, for every ordinal $\lambda$, the principle $AC_{\lambda}$ holds: if $\neq\varnothing, Y\varnothing\notin Y$, and $|Y|\leq|\lambda|$, then there exists a function $f$ with domain $Y$ such that $f(y)\in y$ for all $y\in Y$.
This axiom should be distinguished from the trivial principle asserting choice for families already equipped with specified well–orderings, which is provable in $\mathsf{ZF}$. The present form asserts the existence of a uniform choice function for arbitrary well–orderable families, without any fixed well–orderings given in advance.
As shown by Pincus (see [7]), $\mathsf{AC}_{\mathsf{WO}}$ is equivalent to the statement that $\aleph(X)=\aleph^{*}(X)$ for all sets $X$. This is the form of $\mathsf{AC}_{\mathsf{WO}}$ used implicitly in Ryan–Smith’s localization theorem [17].
@@ 2 Preliminaries and black-box iteration technology
    Set-length and class-length versions of symmetric iterations (cf. [5, §2.1]).
@@
    Definitions and basic properties of countable-support symmetric iterations (cf. [5, Def. 3.1]).
@@
    A $\mathsf{DC}$-preservation theorem for the relevant support (cf. [5, Thm. 3.20]).
@@
    Generation and completeness/normality of limit filters in the relevant cofinality cases (cf. [5, Thm. 5.5]).
@@
    A characterization of HS names by “small support” assuming the stage model satisfies $\mathsf{DC}$ (cf. [5, Lem. 3.14]).
In our application, every stage model satisfies $\mathsf{DC}$ (Proposition 5.65), so the support characterization may be used freely throughout the iteration.
@@
Ryan–Smith localization ingredients (used later). We will repeatedly appeal to two results from [17], which we record in Section 4 in the exact forms used in this paper:
  • Injective upgrade: for every set $S$, $\mathsf{SVC}(S)$ implies $\mathsf{SVC}^{+}(\mathcal{P}(S))$ (Fact 4.3).
  • Localization of $\mathsf{PP}$: assuming $\mathsf{SVC}^{+}(T)$, one has
    $\mathsf{PP}\ \Longleftrightarrow\ \bigl(\mathsf{PP}\!\restriction T\wedge\mathsf{AC}_{\mathsf{WO}}\bigr)$
    (Theorem 4.5, i.e. [17, Proposition 3.17]).
Accordingly, our strategy is to start from a Cohen symmetric seed $S$ with $\mathsf{SVC}(S)$, take $T:=\mathcal{P}(S)$, force $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ while preserving $\mathsf{SVC}(S)$, and then invoke Theorem 4.5 to conclude global $\mathsf{PP}$.
@@ 3.1 The forcing, group action, and the countable-support filter
But $\pi_{i}\operatorname{Fix}(E_{i})\pi_{i}^{-1}=\operatorname{Fix}(\pi_{i}[E_{i}])$, and finite intersections of $\operatorname{Fix}(\cdot)$’s are $\operatorname{Fix}(\bigcup\cdot)$. Let $E:=\bigcup_{i<k}\pi_{i}[E_{i}]$ (finite union of countable sets is countable). Then $\operatorname{Fix}(E)\leq H$. ∎
@@
By Lemma 3.4 (in $V$) choose $E_{n}\in[\omega_{1}]^{\leq\omega}$ with $\operatorname{Fix}(E_{n})\leq H_{n}$. Let $E:=\bigcup_{n<\omega}E_{n}$, which is countable in $V$ since $V\models\mathsf{ZFC}$. Then $\operatorname{Fix}(E)=\bigcap_{n<\omega}\operatorname{Fix}(E_{n})\leq\bigcap_{n<\omega}H_{n}$, hence $\bigcap_{n<\omega}H_{n}\in\mathscr{F}$. ∎
@@ 3.2 CCC and dependent choice
Remark 3.10 (Metatheoretic countability vs. internal countability).
Throughout the construction of HS names, “countable” means countable in the ground model $V$ (e.g. $E\in[\omega_{1}]^{\leq\omega}$ is evaluated in $V$). This notion is used only to ensure that $\operatorname{Fix}(E)\in\mathscr{F}$ (Definition 3.3), hence that the relevant names are symmetric. In particular, when we form supports such as $E=\bigcup_{n<\omega}E_{n}$ in the metatheory, we use that $V\models\mathsf{ZFC}$, so a countable union of countable sets (in $V$) is countable (in $V$). We do not claim that these supports are countable in the symmetric model $\mathcal{N}$.
@@
Let $D_{n+1}:=\bigcup_{p\in D_{n}}D_{n+1,p}$. Since $D_{n}$ is an antichain, if $q\in D_{n+1,p}$ and $q^{\prime}\in D_{n+1,p^{\prime}}$ with $p\neq p^{\prime}$, then $q\perp q^{\prime}$ (otherwise $p$ and $p^{\prime}$ would be compatible). Hence $D_{n+1}$ is an antichain. Moreover, working in the metatheory $V\models\mathsf{ZFC}$ (Remark 3.10), $D_{n+1}$ is countable because it is a countable union (indexed by the countable set $D_{n}$) of countable sets $D_{n+1,p}$. Finally, $D_{n+1}$ is a maximal antichain below $p_{\ast}$: given any $r\leq p_{\ast}$, pick $p\in D_{n}$ compatible with $r$ (maximality of $D_{n}$), let $r^{\prime}\leq r,p$, and then pick $q\in D_{n+1,p}$ compatible with $r^{\prime}$ (maximality of $D_{n+1,p}$).
@@
a countable union of $V$-countable subsets of $\omega_{1}$, hence $V$-countable (Remark 3.10). Therefore $\operatorname{Fix}(E_{n})\in\mathscr{F}$ and $\operatorname{Fix}(E_{n})\leq\operatorname{Stab}(\dot{a}_{n})$. Then every $\pi\in\operatorname{Fix}(E_{n})$ fixes each $p\in D_{n}$, fixes $\dot{A}$ and $\dot{R}$, and fixes each $\dot{a}_{n,p}$. Therefore $\pi\dot{a}_{n}=\dot{a}_{n}$, i.e. $\operatorname{Fix}(E_{n})\leq\operatorname{Stab}(\dot{a}_{n})$, so $\dot{a}_{n}$ is symmetric; hereditariness follows since the names appearing in $\dot{a}_{n}$ are among the $\dot{a}_{n,p}\in\operatorname{HS}$. Thus $\dot{a}_{n}\in\operatorname{HS}$ for all $n$.
Let $E:=\bigcup_{n<\omega}E_{n}$. Then $E$ is $V$-countable (Remark 3.10), hence $\operatorname{Fix}(E)\in\mathscr{F}$, and define the sequence name
@@ 3.3 Countable supports for $\operatorname{HS}$-names
Remark 3.17 (Symmetry support vs. forcing support).
The condition $\operatorname{Fix}(E)\leq\operatorname{Stab}(\tau)$ ($\tau$ has symmetry support EE”) means that $\tau$ is invariant under permutations fixing $E$ pointwise. This does not imply that $\tau$ is a $\mathbb{P}_{E}$–name, nor that $\tau^{G}$ lies in $V[G_{E}]$.
Indeed, letting $\dot{A}=\{(\dot{c}_{\alpha},\mathbf{1}):\alpha<\omega_{1}\}$ be the canonical name for the Cohen set $A=\{c_{\alpha}:\alpha<\omega_{1}\}$, we have $\operatorname{Stab}(\dot{A})=\operatorname{Sym}(\omega_{1})$, so $\dot{A}$ has symmetry support $\emptyset$. Nevertheless, $\dot{A}$ is not a $\mathbb{P}_{E}$–name for any countable $E$, and $A\notin V[G_{E}]$ for every countable $E\subseteq\omega_{1}$.
Lemma 3.18 (Subextension evaluation for $\mathbb{P}_{E}$-names).
Let $E\subseteq\omega_{1}$, let $\mathbb{P}_{E}=\operatorname{Add}(\omega,E)\leq\mathbb{P}=\operatorname{Add}(\omega,\omega_{1})$, and let $\tau$ be a $\mathbb{P}_{E}$–name (viewed as a $\mathbb{P}$–name via the inclusion). If $G\subseteq\mathbb{P}$ is VV–generic and $G_{E}:=G\cap\mathbb{P}_{E}$, then
$\tau^{G}=\tau^{G_{E}}\in V[G_{E}].$
@@
By induction on the rank of $\tau$. Using the inclusion $\mathbb{P}_{E}\hookrightarrow\mathbb{P}$, the valuation recursion for $\tau^{G}$ only consults conditions from $G$ that appear in $\tau$, and these conditions all lie in $\mathbb{P}_{E}$, hence belong to $G_{E}$. Applying the inductive hypothesis to subnames of $\tau$ yields $\tau^{G}=\tau^{G_{E}}$. ∎
@@ 3.4 The canonical Cohen set $A$ and failure of choice
Unlike the finite-support generalized Cohen model (which produces a Dedekind-finite $A$; cf. [15, §2.4]), the present countable-support filter allows many countable subsets of $A$: for each countable $E\subseteq\omega_{1}$ in $V$, the name $\dot{A}_{E}:=\{(\dot{c}_{\alpha},\mathbf{1}):\alpha\in E\}$ is hereditarily symmetric, so $A_{E}:=\{c_{\alpha}:\alpha\in E\}\in\mathcal{N}$.
Lemma 3.23 (Coding a countable set of Cohen reals by an element of $A^{\omega}$).
Let $E\in[\omega_{1}]^{\leq\omega}$ be countable in $V$. Fix in $V$ a surjection $e:\omega\twoheadrightarrow E$. Then the sequence
$s_{E}:=\langle c_{e(n)}:n\in\omega\rangle$
is an element of $A^{\omega}$ in $\mathcal{N}$.
Proof.
Each $\dot{c}_{e(n)}$ is hereditarily symmetric with support $\{e(n)\}$, hence the sequence name $\langle\dot{c}_{e(n)}:n\in\omega\rangle$ is hereditarily symmetric with support $E$. Therefore $s_{E}\in\mathcal{N}$, and clearly $s_{E}\in A^{\omega}$. ∎
Lemma 3.24 (A forced nonempty name yields a ground-coded witness).
Let $\dot{X}$ be a $\mathbb{P}$–name and suppose $p\Vdash``\dot{X}\neq\varnothing$". Then there exist $q\leq p$ and a $\mathbb{P}$–name $\dot{x}$ such that $q\Vdash``\dot{x}\in\dot{X}$".
Proof.
This is the standard forcing-theorem witness extraction for $\exists$. From $p\Vdash\exists x\,(x\in\dot{X})$ we obtain $q\leq p$ and a name $\dot{x}$ with $q\Vdash\dot{x}\in\dot{X}$. ∎
@@ 3.5 A safe $\mathsf{SVC}$ seed in $\mathcal{N}$
We use the standard “continuity” analysis of Cohen-type symmetric models. In the basic Cohen model $V(A)$ generated by an unordered set $A$ of Cohen reals, a key lemma (sometimes called a continuity or indiscernibility lemma) shows that statements over $V(A)$ with parameters from $V$ are already decided in a subextension generated by finitely many members of $A$; see, e.g., Felgner [4, p. 133] or Blass [3, Prop. 1.2] (and the discussion in [6, Lem. 2.8]).
More precisely, for each $x\in X$ there is a $\mathbb{P}$–name $\tau_{x}\in V$ and a sequence $s_{x}\in A^{\omega}$ (coding the countable symmetry support relevant to $\tau_{x}$) such that $x$ is recovered uniformly from the pair $(\tau_{x},s_{x})$. Enumerating the set of $\mathbb{P}$–names in $V$ by ordinals yields the required surjection from $A^{\omega}\times\eta$ onto $X$ for some ordinal $\eta$.
The same argument applies to our Cohen symmetric model $\mathcal{N}$ built from $\operatorname{Add}(\omega,\omega_{1})$ with the countable-support symmetry filter: every set $X\in\mathcal{N}$ is obtained (uniformly in $X$) from $V$ and a countable sequence of members of $A$, and therefore is a surjective image of $A^{\omega}\times\eta$ for some ordinal $\eta$. Equivalently, $\mathcal{N}\models\mathsf{SVC}(A^{\omega})$. ∎
@@ 4 Reduction blueprint: global $\mathsf{PP}$ from localized $\mathsf{PP}$
For every set $S$, $\mathsf{SVC}(S)$ implies $\mathsf{SVC}^{+}(\mathcal{P}(S))$. [17, §2.1, Fact]
@@
[17, Proposition 3.17] (stated there with parameter “SS”; we apply it with $S:=T$). As this argument is critical to our goals, we demonstrate a proof here.
Proof.
($\Rightarrow$) If $\mathsf{PP}$ holds, then $\mathsf{PP}\!\restriction T$ is immediate. Moreover, $\mathsf{PP}$ implies $\mathsf{AC}_{\mathsf{WO}}$ (e.g. via the consequence $(\forall X)\,\aleph(X)=\aleph^{\ast}(X)$; see [17, Proposition 3.17 and footnote 4] for references).
($\Leftarrow$) Assume $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$, and let $\pi:Y\twoheadrightarrow X$ be a surjection. We must produce an injection $X\hookrightarrow Y$.
Apply $\mathsf{SVC}^{+}(T)$ to the set
$Z\;:=\;(X\times\{0\})\ \cup\ (Y\times\{1\}),$
obtaining an ordinal $\eta$ and an injection $j:Z\hookrightarrow T\times\eta$. Define injections
$\pi_{X}:X\to T\times\eta,\quad\pi_{X}(x)=j(\langle x,0\rangle),\qquad\pi_{Y}:Y\to T\times\eta,\quad\pi_{Y}(y)=j(\langle y,1\rangle),$
and let
$A:=\operatorname{ran}(\pi_{X})\subseteq T\times\eta,\qquad B:=\operatorname{ran}(\pi_{Y})\subseteq T\times\eta.$
Then $\pi_{X}:X\to A$ and $\pi_{Y}:Y\to B$ are bijections, hence have inverses $\pi_{X}^{-1}:A\to X$ and $\pi_{Y}^{-1}:B\to Y$. Define
$f\;:=\;\pi_{X}\circ\pi\circ\pi_{Y}^{-1}\;:\;B\twoheadrightarrow A.$
It therefore suffices to show that whenever $A,B\subseteq T\times\eta$ and $f:B\twoheadrightarrow A$, there is an injection $g:A\hookrightarrow B$.
Claim 2. (Slicing lemma for $T\times\eta$) Let $\eta$ be an ordinal. Assume $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$. If $A,B\subseteq T\times\eta$ and $f:B\twoheadrightarrow A$, then there is an injection $g:A\hookrightarrow B$.
Proof of Claim 4.
Treat $f$ as a partial map on $T\times\eta$ with domain $B$. For each $\langle t,\alpha\rangle\in A$, define
$\varepsilon_{t,\alpha}\;:=\;\min\bigl\{\varepsilon<\eta:(\exists s\in T)\ \bigl(\langle s,\varepsilon\rangle\in B\ \wedge\ f(\langle s,\varepsilon\rangle)=\langle t,\alpha\rangle\bigr)\bigr\}.$
(Nonempty since $f$ is surjective and $\min$ exists since $\eta$ is an ordinal.)
For $\varepsilon,\alpha<\eta$, define the slices
$A^{\varepsilon}_{\alpha}\;:=\;\{\,t\in T:\langle t,\alpha\rangle\in A\ \wedge\ \varepsilon_{t,\alpha}=\varepsilon\,\}\ \subseteq\ T,$
and
$B^{\varepsilon}_{\alpha}\;:=\;\{\,s\in T:\langle s,\varepsilon\rangle\in B\ \wedge\ f(\langle s,\varepsilon\rangle)=\langle t,\alpha\rangle\text{ for some }t\in A^{\varepsilon}_{\alpha}\,\}\ \subseteq\ T.$
Define $h^{\varepsilon}_{\alpha}:B^{\varepsilon}_{\alpha}\to A^{\varepsilon}_{\alpha}$ by
$h^{\varepsilon}_{\alpha}(s)=t\quad\text{iff}\quad f(\langle s,\varepsilon\rangle)=\langle t,\alpha\rangle.$
This is well-defined (since $f$ is a function), and it is a surjection: if $t\in A^{\varepsilon}_{\alpha}$, then by definition of $\varepsilon_{t,\alpha}=\varepsilon$ there exists $s$ with $\langle s,\varepsilon\rangle\in B$ and $f(\langle s,\varepsilon\rangle)=\langle t,\alpha\rangle$, hence $s\in B^{\varepsilon}_{\alpha}$ and $h^{\varepsilon}_{\alpha}(s)=t$.
By $\mathsf{PP}\!\restriction T$ applied to the surjection $h^{\varepsilon}_{\alpha}:B^{\varepsilon}_{\alpha}\twoheadrightarrow A^{\varepsilon}_{\alpha}$, there exists an injection $i^{\varepsilon}_{\alpha}:A^{\varepsilon}_{\alpha}\hookrightarrow B^{\varepsilon}_{\alpha}$ whenever $A^{\varepsilon}_{\alpha}\neq\varnothing$.
Let
$E\;:=\;\{\,\langle\varepsilon,\alpha\rangle\in\eta\times\eta:A^{\varepsilon}_{\alpha}\neq\varnothing\,\}.$
Then $E$ is well-orderable (e.g. by the lexicographic order on $\eta\times\eta$), so by $\mathsf{AC}_{\mathsf{WO}}$ we may choose a function $c$ with
$c(\varepsilon,\alpha)\in\{\,i:i\text{ is an injection }A^{\varepsilon}_{\alpha}\hookrightarrow B^{\varepsilon}_{\alpha}\,\}\quad\text{for each }\langle\varepsilon,\alpha\rangle\in E.$
Define $g:A\to T\times\eta$ by
$g(\langle t,\alpha\rangle)\;:=\;\bigl\langle\,c(\varepsilon_{t,\alpha},\alpha)(t),\ \varepsilon_{t,\alpha}\,\bigr\rangle.$
Then $g(\langle t,\alpha\rangle)\in B$: indeed, $c(\varepsilon_{t,\alpha},\alpha)(t)\in B^{\varepsilon_{t,\alpha}}_{\alpha}$, so by definition of $B^{\varepsilon_{t,\alpha}}_{\alpha}$ we have $\bigl\langle c(\varepsilon_{t,\alpha},\alpha)(t),\varepsilon_{t,\alpha}\bigr\rangle\in B$.
Finally, $g$ is injective. If $g(\langle t,\alpha\rangle)=g(\langle t^{\prime},\alpha^{\prime}\rangle)$, then $\varepsilon_{t,\alpha}=\varepsilon_{t^{\prime},\alpha^{\prime}}=:\varepsilon$ from equality of second coordinates. Let $s:=c(\varepsilon,\alpha)(t)=c(\varepsilon,\alpha^{\prime})(t^{\prime})$ be the common first coordinate. Then $s\in B^{\varepsilon}_{\alpha}\cap B^{\varepsilon}_{\alpha^{\prime}}$. But if $\alpha\neq\alpha^{\prime}$ this is impossible: $s\in B^{\varepsilon}_{\alpha}$ implies $f(\langle s,\varepsilon\rangle)$ has second coordinate $\alpha$, while $s\in B^{\varepsilon}_{\alpha^{\prime}}$ implies the second coordinate is $\alpha^{\prime}$, contradicting that $f$ is a function. Hence $\alpha=\alpha^{\prime}$, and then $c(\varepsilon,\alpha)$ is injective, so $t=t^{\prime}$. Thus $\langle t,\alpha\rangle=\langle t^{\prime},\alpha^{\prime}\rangle$. ∎
By Claim 4 there is an injection $g:A\hookrightarrow B$. Then
$i\;:=\;\pi_{Y}^{-1}\circ g\circ\pi_{X}\;:\;X\to Y$
is an injection, as required. Therefore $\mathsf{PP}$ holds. ∎
@@
This is witnessed by Cohen’s first model; see [17, Proposition 3.22]. ∎
@@
This is witnessed by the Feferman-style model $N_{\aleph_{1}}$ of Truss; see [17, Proposition 3.23]. ∎
@@
The preceding propositions are included only to justify (conceptually) why the localization theorem is stated with both components present: the “choice” side $\mathsf{AC}_{\mathsf{WO}}$ and the “localized comparability” side $\mathsf{PP}\!\restriction T$. We will force both conjuncts $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ in Theorem 4.5. In Ryan–Smith’s analysis the corresponding reduction is sharp: in general one cannot drop the displayed conjuncts (nor substantially weaken the hypotheses) without losing the equivalence; see [17, Props. 3.22 and 3.23].